OpenStudy (anonymous):

Write the equation in a standard form ax^2+bx+c=0 , then give the values of a,b, and c. (2x-1)(3x+2)=2(2x^2+3)

OpenStudy (lalaly):

\[\large{6x^2+4x-3x-2}=4x^2+6\]

OpenStudy (lalaly):

\[\large{6x^2-4x^2+4x-3x-2-6=0}\]

OpenStudy (lalaly):

\[\large{2x^2+x-8=0}\]

OpenStudy (lalaly):

a=2 b=1 c=-8

OpenStudy (anonymous):

i get the same:)

OpenStudy (lalaly):

:D